mathslover (mathslover):

If a, b and c are positive real numbers such that a + b + c = 3 then prove that : \(\sqrt{a^3 + 3b} + \sqrt{b^3+ 3c} +\sqrt{c^3 +3a} \ge 6\)

OpenStudy (mathmale):

Hello, Kush, Great to hear from you! Regarding the math problem you've posted: I don't immediately see the way to a complete proof. I do have a couple of suggestions to start with, however: 1) Experiment with actual values of a, b and c such that a+b+c=3. It goes without saying that a=1, b=1, c=1 is one such possibility. Substitute these three values back into the original equation. Is the resulting equation true or false? What, if anything, can you learn from this? 2) Experiment with eliminating one variable: for example, c-3-a-b. 3) Try eliminating the square roots by squaring both sides of the equation. You could, for example, take the original equation, move one of the three square roots to ther right side, and then square each of the (binomial) sides of the resulting equation. This would produce another equation that still has one or more square root operators, but you could eliminate it / them in the same manner. Best to you!

mathslover (mathslover):

Sir, thanks a lot for those suggestions. 1) for a = b = c = 1, I get that original inequality true with equality sign. That is, the LHS equals 6. The resulting equation is true. I'm not sure if I will be able to learn anything from this. What I think is that a , b and c should be greater than or equal to 1 for the inequality to hold. 2) Eliminating a variable (taking your example into consideration) gives : \(c = 3 - (a+b)\) \(\sqrt{a^3 + 3b} + \sqrt{b^3 + 3(3 - (a+b))} + \sqrt{(3-(a+b))^3 + 3a} \ge 6 \\ \sqrt{a^3+ 3b} + \sqrt{b^3 + 9 - 3a - 3b} +\sqrt{(3-(a+b))^3 + 3a} \ge 6\) While if I keep simplifying it I don't get anything useful. Should I solve it further? 3) \(\sqrt{a^3 + 3b} + \sqrt{b^3 + 3c} \ge 6 - \sqrt{c^3 + 3a} \\ \text{Squaring Both Sides} \\ a^3 + 3b + b^3 + 3c + 2\sqrt{(a^3 + 3b)(b^3 + 3)} \ge 36 + c^3 + 3a - 12\sqrt{c^3 + 3a} \)

mathslover (mathslover):

I tried to follow your guidelines, but there is nothing satisfactory yet. Though, I usually don't give up so earlier, but, I'm not finding any proper way to go on from the equations I got. As the square roots will be irritating later. \((a^3 + b^3 + 3(b +c) ) + 2\sqrt{(a^3 + 3b)(b^3 +3c)} \ge 36 + c^3 + 3a - 12\sqrt{c^3 +3a} \\ 2\sqrt{(a^3 + 3b)(b^3 +3c)} \ge 36 + c^3 - (a^3 + b^3 + 3(b+c)) + 3a - 12\sqrt{c^3 + 3a} \\ 2\sqrt{(a^3 + 3b)(b^3 +3c)} + 12\sqrt{c^3 + 3a} \ge 36 + c^3 - a^3 - b^3 - 3(b+c) \) Again, if I will square both sides, it will become more complex. :(

OpenStudy (mathmale):

Challenging problem, yes. As you can see, I've approached it by identifying and considering a variety of approaches, since I've not encountered such a proof problem before. At this point I'd suggest you try brainstorming on other possible approaches and then try them out. I understand you live in a relatively isolated place, but am wondering whether you have a colleague (fellow student) with whom you could discuss this and other problems. Kush, i need to get off the 'Net soon. I hope to exchange further messages with you very soon. Best to you.

mathslover (mathslover):

I will try to discuss this with my brother. Sure Sir. I will let you know if I get any method to solve this. Again, thanks a lot for your responses.

OpenStudy (experimentx):

somehow i got the weaker bound for it ...it's equal to \( \sqrt {12} \)

mathslover (mathslover):

What is equal to \(\sqrt{12}\) ? The LHS?

OpenStudy (experimentx):

RHS.

mathslover (mathslover):

Oh, wait, which equation are you referring to?

OpenStudy (experimentx):

\[ \sqrt{a^3 + 3b} + \sqrt{b^3+ 3c} +\sqrt{c^3 +3a} \ge \sqrt{12} \]

mathslover (mathslover):

Oh, okay, so you have proved this for \(\ge 2\sqrt{3}\) while we had to prove for \(\ge 6\) . Though, this doesn't give us the required proof, but I will love to see how you proved that,

OpenStudy (experimentx):

use the fact that \[ a+b+c \ge \sqrt{a^2 + b^2 + c^2}\] and \[ a^3 + b^3+c^3 \ge 3 , a+b+c=3\]

OpenStudy (experimentx):

there is long way to prove such inequalities using calculus, it called the method of lagrange multiplies ... but it's not nice and way too long.

mathslover (mathslover):

\(3 \ge \sqrt{a^2 + b^2 + c^2}\) and \(a^3 + b^3 + c^3 \ge \sqrt{a^2 + b^2 + c^2}\) I have not studied Lagranges' Multiples yet, and if it is way too long, then let us prefer to find a shorter method.

OpenStudy (experimentx):

no ... make subsitution \[ \sum \sqrt{a^3 + 3b} \ge \sqrt{\sum a^3 + 3b }\]

mathslover (mathslover):

Okay, but, how did you get to that substitution ?

OpenStudy (experimentx):

let x =a^3+3b

OpenStudy (experimentx):

this is much more weaker than the inequality you state, it's useless to do it.

mathslover (mathslover):

Yeah..

mathslover (mathslover):

\( 9 = a^2 + b^2 + c^2 + 2(ab + bc + ca) \ \ a^2 + b^2 + c^2 = 9 - 2(ab + bc + ca) \\ a^3 + b^3 +c^3 \ge \sqrt{a^2 + b^2 + c^2}\\ \implies a^3 + b^3 + c^3 \ge \sqrt{9 - 2(ab + bc + ca) } \\ \implies (a^6 + b^6 + c^6 + 2(a^3 b^3 + b^3c^3 + c^3a^3) \ge 9 -2ab -2bc - 2ac \\ \implies a^6 + b^6 + c^6 + 2ab(a^2b^2 + 1) + 2bc(b^2c^2 +1) + 2ac(c^2a^2 +1) \ge 9\) While, I'm thinking to do something with this ^ :P not sure whether this will give us anything useful or not.

mathslover (mathslover):

While at the other part of my mind, I am thinking to prove this : \(\sqrt{a^3 + 3b} \ge 2\) ; \(\sqrt{b^3 +3c} \ge 2\) ; \(\sqrt{c^3 +3a} \ge 2\) This will be much easier for us. Let us just take one equation into consideration. \(\boxed{\sqrt{a^3 +3b} \ge 2}\) \(a^3 +3b \ge 4\) This should be true.. hmm!

OpenStudy (experimentx):

no that's incorrect ..

OpenStudy (vishweshshrimali5):

Try using Cauchy Schwarz inequality

OpenStudy (experimentx):

just choose a=b=0, c=3

mathslover (mathslover):

Oh.. okay, yep, I used wrong method.

mathslover (mathslover):

@vishweshshrimali5 - can you explain it a litle bit, like how will it help us?

OpenStudy (vishweshshrimali5):

First of all do you know about it ?

OpenStudy (vishweshshrimali5):

http://en.wikipedia.org/wiki/Cauchy%E2%80%93Schwarz_inequality

OpenStudy (experimentx):

it seemed like you can find upper bound using cauchy swarz.

mathslover (mathslover):

https://fbcdn-sphotos-a-a.akamaihd.net/hphotos-ak-xpa1/t1.0-9/10371689_10152501448231552_8766121181245523511_n.jpg Here is from where I got this identity. I don't know what has the user done there, whether he is proving this identity there or is doing something else.

OpenStudy (vishweshshrimali5):

He has used CS theorem.

OpenStudy (vishweshshrimali5):

Cauchy Schwarz inequality = CS ineq.

mathslover (mathslover):

Oh, so he has proved it somewhere ? :P

mathslover (mathslover):

Well, to be honest, I don't understand the CS inequality. Is this the only way to prove this identity? Well, I'm still trying to understand it...

OpenStudy (vishweshshrimali5):

It basically says that: \[\large{(xa+yb)^2 \le (x^2 + y^2)(a^2 + b^2)}\]

OpenStudy (vishweshshrimali5):

As for the proof, search on internet.

OpenStudy (vishweshshrimali5):

No he has not proved this ineq. anywhere, it is a very common ineq. among olympiad students. So, he has taken it that the student knows it.

OpenStudy (vishweshshrimali5):

This will give you the very basic idea of the ineq.: \(\color{blue}{\text{Originally Posted by}}\) @vishweshshrimali5 It basically says that: \[\large{(xa+yb)^2 \le (x^2 + y^2)(a^2 + b^2)}\] \(\color{blue}{\text{End of Quote}}\)

mathslover (mathslover):

Oh, I see. I will try hard to understand this and get through the problem. Will post the solution I get. @experimentX , @mathmale and @vishweshshrimali5 - You guys are awesome. Thanks a lot Everyone.

OpenStudy (vishweshshrimali5):

np at all. And others did the most work, I just gave one single hint. :)

OpenStudy (experimentx):

http://www.aam.org.in/site/st_material/14.pdf

mathslover (mathslover):

@experimentX - That pdf is really nice. I would love to read it :-) Thanks.